Calculate the mutual inductance between two coils that are inductively coupled if 112V is induced in the secondary when the 60-cylce primary current is 2.8 Amp.


1
Expert's answer
2021-05-24T11:42:01-0400

"i=i_0\sin\omega t"


"E_2=-M_{21}\frac{di}{dt}=-M_{21}i_0\omega\cdot\cos\omega t" . So,


"M_{21}=E_2/(i_0\omega)=112/(2.3\cdot2\cdot3.14\cdot60)=0.297\ (H)"